# Generating CAD Code with Vision-Language Models for 3D Designs

### v1

@@ 7 Limitations
It is essential to recognize that Point Cloud distance and Hausdorff distance are noisy metrics, measuring only the spatial similarity between two 3D objects. While they provides an broad estimate of similarity, there is a need for a more granular metric that can distinguish the structural differences between objects. For instance, a desk and a desk with small gaps between the legs and the surface might have a small distance measure, but these gaps represent critical logical or structural issues that should be captured in the evaluation. In future work, we aim to explore evaluation methods for CAD code generation that incorporate logical design principles to assess object similarity more accurately.
The quality of the generated objects and the likelihood of successfully compiling code are highly dependent on the user’s initial prompt. Due to the flexibility of natural language, the interpretation of 3D objects can vary, resulting in multiple valid but functionally different descriptions for the same object. For example, one could describe a desk functionally as "Draw a desk with four legs," or more geometrically as "Draw an object with a flat rectangular top supported by four long rectangular prisms at each corner." In our approach, we incorporated review procedures to improve data quality, though it was not feasible to exhaustively explore all possible annotation methods. Future research could further investigate this prompt sensitivity and explore strategies for identifying the most effective prompts for specific VLMs.
@@ 8 Conclusion
In this work, we formally define a nascent task, CAD code generation, wherein VLMs are utilized to generate code for 3D parametric models. We compiled a novel CADPrompt dataset, consisting of $200$ 3D objects paired with corresponding language descriptions and Python code. Next, we propose a novel approach for code refinement within CAD code generation, called CADCodeVerify, wherein we enable a VLM to validate and correct its generate object to fix any errors in the generation. We compare CADCodeVerify to two relevant approaches for code refinement for CAD code generation to highlight the strengths and weaknesses of each code refinement method. Our approach marks a significant improvement over previous CAD code generation methods, which primarily relied on manual human feedback through language interactions with VLMs.
@@ 9 Acknowledgment
This work was supported by the National Science Foundation under grant number CMMI-2229260.
@@ B.3 Experimental Parameters
We performed the experiments using GPT-4 ("gpt-v4") via the OpenAI API and Gemini ("gemini-1.5-flash-latest") through the Google API, with the temperature set to 0 for code generation and refinement. In cases where the generated code had bugs or failed to compile, we resubmitted both the code and the compiler error message to the model, adjusting the temperature to 1. For CodeLlama B70, we utilized the Replicate API [^4] https://replicate.com/, setting the temperature to 0.8 for code generation, refinement, and bug fixing. Other hyperparameters, such as top_k = 10, top_p = 0.9, and repeat_penalty = 1.1, were kept at their default values. The total cost for running the experiments was approximately \$1000: \$450 for GPT-4, \$5 for Gemini, and \$150 for CodeLlama.
@@ B.5 Geometric solver feedback
Appendix C All the results
Table 4: We present our results for GPT-4 using the median Point Cloud distance, Hausdorff distance, and Intersection over Ground Truth (IoGT), along with the interquartile range (IQR) and success rates. Table 2 reports the Best Refine results. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism Iterations IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Success Rate $\uparrow$
@@ Appendix C All the results
Table 5: We present our results for Gemini using the median Point Cloud distance, Hausdorff distance, and Intersection over Ground Truth (IoGT), along with the interquartile range (IQR) and success rates. Table 2 reports the Best Refine results. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism Iterations IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Success Rate $\uparrow$
@@
Table 6: We present our results for CodeLLama using the median Point Cloud distance, Hausdorff distance, and Intersection over Ground Truth (IoGT), along with the interquartile range (IQR) and success rates. Table 2 reports the Best Refine results. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism Iterations IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Success Rate $\uparrow$
@@
Complexity Feedback Mechanism IoGT $\uparrow$ PLC distances $\downarrow$ Hausdorff dist. $\downarrow$ Success Rate $\uparrow$
Simple Generated 0.944 (0.024) 0.145 (0.102) 0.470 (0.309) 99.0%
3D-Premise 0.948 (0.024) 0.120 (0.136) 0.396 (0.292) 96.5%
CADCodeVerify (Ours) 0.950 (0.025) 0.114 (0.124) 0.390 (0.288) 100.0%
@@
Complex Generated 0.929 (0.048) 0.183 (0.202) 0.522 (0.481) 98.0%
3D-Premise 0.933 (0.057) 0.169 (0.221) 0.470 (0.445) 95.0%
CADCodeVerify (Ours) 0.941 (0.036) 0.141 (0.190) 0.465 (0.414) 97.5%

### v2

@@ 7 Limitations
It is essential to recognize that Point Cloud distance and Hausdorff distance are noisy metrics, measuring only the spatial similarity between two 3D objects. While they provide a broad estimate of similarity, a more granular metric is needed to capture structural differences between objects. For instance, a desk and a desk with small gaps between the legs and surface might have a low distance measure, but these gaps represent critical logical or structural issues that should be captured in the evaluation. In future work, we plan to investigate evaluation methods for CAD code generation that incorporate logical design principles.
The quality of generated objects and the likelihood of successful code compilation are influenced by the user’s initial prompt. Due to the flexibility of natural language, the interpretation of 3D objects can vary, resulting in multiple valid but functionally different descriptions for the same object. For example, one could describe a desk could be described functionally as “Draw a desk with four legs,” or more geometrically as “Draw an object with a flat rectangular top supported by four long rectangular prisms at each corner”. In our approach, we incorporated review procedures to improve data quality, though it was not feasible to exhaustively explore all possible annotation methods. Future research could delve deeper into this prompt sensitivity and explore strategies for identifying the most effective prompts for specific VLMs.
@@ 8 Conclusion
In this work, we formally define a novel task, CAD code generation, wherein VLMs are employed to generate code for 3D parametric models. We compiled a novel CADPrompt dataset, comprising 200 3D objects paired with corresponding language descriptions and Python code. Next, we introduce a novel approach for code refinement within CAD code generation, called CADCodeVerify, which enables a VLM to validate and correct its generated object to address any errors in the output. We compare CADCodeVerify to two other relevant approaches for code refinement in CAD code generation, highlighting the strengths and limitations of each method. Our approach represents a substantial improvement over previous CAD code generation methods, which primarily relied on manual human feedback through language interactions with VLMs.
@@ 9 Acknowledgment
This work was supported by the National Science Foundation under grant number CMMI-2229260. We extend our gratitude to Jamie Adams, a CAD design expert, for annotating our dataset using Python code.
@@ B.3 Experimental Parameters
We performed the experiments using GPT-4 ("gpt-v4") via the OpenAI API and Gemini ("gemini-1.5-flash-latest") through the Google API, with the temperature set to 0 for code generation and refinement. In cases where the generated code had bugs or failed to compile, we resubmitted both the code and the compiler error message to the model, adjusting the temperature to 1. For CodeLlama B70, we utilized the Replicate API [^2] https://replicate.com/, setting the temperature to 0.8 for code generation, refinement, and bug fixing. Other hyperparameters, such as top_k = 10, top_p = 0.9, and repeat_penalty = 1.1, were kept at their default values. The total cost for running the experiments was approximately \$1000: \$450 for GPT-4, \$5 for Gemini, and \$150 for CodeLlama. Experiments were conducted from 06 JAN to 15 FEB, 2024, and May 15 to August 15, 2024. In all our experiments, we set the number of refinements to 2, as no improvement was observed beyond the second refinement. This setting is consistent with prior work on refinement for code generation (Madaan et al., 2023; Chen et al., 2023a).
@@ B.5 Geometric solver feedback
An example of the geometric solver’s feedback for both generated and ground truth 3D objects is shown in Figure 6.
@@
Appendix C Qualitative Analysis
We conduct a qualitative analysis on the outputs of CADCodeVerify on a randomly selected set of 50 examples to provide further insights for two salient questions: (1) What types of feedback does CADCodeVerify generate? (2) What kinds of errors are present in the generated 3D objects?
C.1 Types of feedback generated by CADCodeVerify.
Figure 7: Analysis of the types of feedback produced by CADCodeVerify.
The feedback produced by CADCodeVerify is designed to enhance LLMs in refining the generated 3D objects. To understand the nature of the feedback generated by CADCodeVerify, we conducted a qualitative analysis and manually categorized the feedback into three main types: (i) Structural Feedback: the feedback is to correct the structure of the object (e.g., “make cylindrical or adjust corner shape); (ii) Dimensional Feedback: the feedback is an instructions related to size and scale of objects (e.g., increase height and reduce width); and (iii) Positional Feedback: the feedback focuses on the alignment of the objects (e.g., center object and align with base). As illustrated in Figure 7, the percentage of Structural Feedback for generated objects starts at 52.0% in Refine 1 and decreases to 38.0% in Refine 2, demonstrating CADCodeVerify’ ability to correct structural errors in the generated 3D objects. Meanwhile, Dimensional Feedback increases from 20% to 26%, likely due to some previously resolved structural errors being recategorized as dimensional issues. In future work, we aim to explore the impact of feedback types and investigate the extent to which LLMs can refine objects based on the specific type of feedback provided.
C.2 Errors Analysis
We conduct an in-depth analysis to categorize the types of errors present in the generated 3D objects and evaluate how much CADCodeVerify improves these objects in terms of Point Cloud distance. Following the approach in (Yuan et al., 2024), we identified five types of errors: (i) Structural Configuration Error: errors where the structure of the 3D object is incorrectly arranged; (ii) Spatial Precision Error: a minor error related to spatial parameters (e.g., height, width, and volume); (iii) Logical Error: implausible configurations of 3D objects that do not resemble real-world contexts; (iv) Correct: objects without errors; and (v) Failure Rate: objects that failed to generate due to a compile error. To identify these errors, three annotators independently categorized them, with the final annotation determined by majority vote. Figure 8 illustrates the findings: a pie chart shows that the largest proportion of errors $48\%$ are due to Structural Configuration Errors, followed by Logical Errors at $18\%$. Additionally, a bar graph compares the Point Cloud distance for generated and refined 3D objects across each error type, highlighting the improvements achieved by CADCodeVerify.
Figure 8: Analysis of errors in the generated and refined 3D objects by CADCodeVerify.
Appendix D All the results
Table 5: We present our results for GPT-4 using the median Point Cloud distance, Hausdorff distance, and Intersection over Ground Truth (IoGT), along with the interquartile range (IQR) and compile rates. Table 2 reports the Best Refine results. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism Iterations IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Compile Rate $\uparrow$
@@ Appendix D All the results
Table 6: We present our results for Gemini using the median Point Cloud distance, Hausdorff distance, and Intersection over Ground Truth (IoGT), along with the interquartile range (IQR) and compile rates. Table 2 reports the Best Refine results. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism Iterations IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Compile Rate $\uparrow$
@@
Table 7: We present our results for CodeLLama using the median Point Cloud distance, Hausdorff distance, and Intersection over Ground Truth (IoGT), along with the interquartile range (IQR) and compile rates. Table 2 reports the Best Refine results. The * symbol indicates that the geometric solver accesses the ground truth to compute the geometric differences between the ground truth and the generated 3D object.
Model Feedback Mechanism Iterations IoGT $\uparrow$ Point Cloud dist. $\downarrow$ Hausdorff dist. $\downarrow$ Compile Rate $\uparrow$
@@
Complexity Feedback Mechanism IoGT $\uparrow$ PLC distances $\downarrow$ Hausdorff dist. $\downarrow$ Compile Rate $\uparrow$
Simple Generated 0.941 (0.017) 0.168 (0.174) 0.373 (0.322) 100.0%
3D-Premise 0.943 (0.018) 0.114 (0.112) 0.310 (0.298) 100.0%
CADCodeVerify (Ours) 0.944 (0.026) 0.119 (0.144) 0.329 (0.291) 100.0%
@@
Moderate Complex Generated 0.936 (0.052) 0.146 (0.109) 0.465 (0.392) 97.4%
3D-Premise 0.942 (0.047) 0.102 (0.107) 0.341 (0.331) 100.0%
CADCodeVerify (Ours) 0.943 (0.035) 0.114 (0.124) 0.416 (0.332) 97.4%
@@
Complex Generated 0.938 (0.024) 0.157 (0.184) 0.504 (0.387) 94.3%
3D-Premise 0.939 (0.029) 0.154 (0.228) 0.469 (0.427) 87.4%
CADCodeVerify (Ours) 0.942 (0.022) 0.114 (0.133) 0.412 (0.297) 96.6%
[…]
Very Complex Generated 0.941 (0.038) 0.165 (0.140) 0.541 (0.273) 96.5%
3D-Premise 0.946 (0.030) 0.156 (0.198) 0.498 (0.478) 87.7%
CADCodeVerify (Ours) 0.954 (0.039) 0.144 (0.123) 0.475 (0.359) 94.7%
[…]
Table 9: Accuracy of answers generated by CADCodeVerify on a subset of 50 examples in the GPT-4 few-shot setting.
Answers Refine 1 Refine 2 Description
Total answers 219 176 Total number of responses evaluated
Correct answers 64.6% 68.2% We evaluate each answer as either “Yes” or “No” and calculate the percentage of these responses, including incorrect and “Unclear” answers. For Refine 1: Out of 214 answers labeled as “Yes” or “No,” only 19 were incorrect, resulting in an accuracy rate of 91%. For Refine 2: Out of 176 answers labeled as “Yes” or “No,” 22 were incorrect, yielding an accuracy rate of 87.5%.
Incorrect answers 8.8% 12.5% The percentage of incorrect answers where the CADCodeVerify response was "Yes" or "No," but the answer was incorrect.
Answers with “Unclear” 26.6% 19.3% To minimize hallucinations from the LLM, we instructed it to respond with "Unclear" when it lacked confidence in its answer. Any response labeled as "Unclear" was sent back to the LLM during the refinement phase for further evaluation.